If f(x)=f(x-2)+x, and f(7)=11, find f(5).
Answer by ikleyn(53765)

.

Take x= 5 and apply the given data.   You will have


    f(7) = f(5) + 7,

or

    11   = f(5) + 7.


It implies


    f(5) = 11 - 7 = 4.    ANSWER

Solved.